# Hannes Alfvéns väg 12, SE-106 91 Stockholm, Sweden

### v1

@@ Abstract
Scalar fields with masses protected by global shift symmetries, commonly referred to as axions, are abundantly used in effective field theories in cosmology and particle physics. However, global symmetries cannot be expected to be protected at the fundamental level. Finding consistent ultra-violet completions for axions is therefore a necessity. In this work, we identify the axion with the position mode of a charged 3-brane in (4+1)-dimensions. The shift symmetry of the axion is then a residual diffeomorphism in the fifth dimension orthogonal to the brane. Meanwhile, the brane is coupled to a flux in the fifth dimension. From the (3+1)-dimensional perspective, this construction generates a mass for the axion and matches previously known proposals in the literature based on the coupling between the axion and a three-form gauge field. In a second step, we uplift this (4+1)-dimensional model to M-theory, where the same three-form is found to couple to the membrane with a (2+1)-dimensional worldvolume. In particular, our proposal also elucidates the duality between the axion and a two-form gauge field in the literature. We show that this dual two-form couples to the boundary of an open membrane in M-theory. Finally, we comment on the relations to and differences from other closed and open string axion monodromy models.
@@ 2.3 Mass Generation from Three-Form Potential
$\int_{\mathcal{M}_{4}}C^{(4)}=\frac{1}{4!}\int_{\mathcal{M}_{4}}\text{d}^{4}x\,\epsilon^{\mu\nu\rho\sigma}\,\partial_{\mu}f^{\text{A}}\,\partial_{\nu}f^{\text{B}}\,\partial_{\rho}f^{\text{C}}\,\partial_{\sigma}f^{\text{D}}\,C_{\mathrm{ABCD}}\,.$ (21)
@@ 2.4 Dimensional Reduction and Anti-Brane
Figure 3: An illustration of the two-brane configuration. Our universe as a 3-brane sits at $y=0$ and its anti-brane at $y=\pi\,R$ , with $y$ the compactified dimension that is orthogonal to the brane. These two branes carry the same charge $\mathcal{Q}$ but with opposite signs. The background value of the four-form field strength $\bar{H}_{01234}$ is a discontinuous function in $y$, which is equal to $q_{\text{S}}$ in the southern semi-circle and $q_{\text{N}}$ in the northern semi-circle.
Here we provide a more self-contained setting that takes into account the back-reaction of the brane charges on the bulk fluxes. As a result, we will show that the axion potential in the four-dimensional EFT becomes multi-valued, similar to models of axion monodromy. The self-consistency of the boundary condition in the compact direction requires that we generalize the action (19) to include an anti 3-brane with opposite charge. For now, we ignore gravitational effects but will get back to the stability of our setting later. We also assume that they have the same tension $\mathcal{T}$ and are placed at $y_{1}=0$ and $y_{2}=R\,\pi$ in the compact dimension. The five-dimensional action (19) is now generalized to be
@@
$\partial_{y}\bar{H}_{01234}=\Bigl{[}\mathcal{Q}_{1}\,\delta\bigl{(}y-y_{1}\bigr{)}+\mathcal{Q}_{2}\,\delta\bigl{(}y-y_{2}\bigr{)}\Bigr{]}\,.$ (34)
We integrate Eq. (34) to obtain
@@
In $S^{(2)}_{\text{rest}}$ , the first term does not make any contribution to the dynamics of the axions and we thus drop it in the following discussion. In the second line, we encounter a term linear in $a_{1}-a_{2}$ . The coefficient of this latter linear terms is proportional to $\partial_{y}\bar{C}_{0123}=\bar{H}_{01234}$ , which is an undetermined constant on the brane itself. Following the rest of the derivations in Section 2.3 by dualizing $\delta C_{\mu\nu\rho}$ , we find the analog of Eq. (30), with
@@
which further implies that the axion theory (39) admits a discrete gauge symmetry,
@@
Here we give some preliminary remarks on the stability of this brane-anti-brane system.  [^10] See related discussions in Burgess:2001fx . Schematically, the gravitational potential $V_{\mathrm{gravity}}$ in the fifth dimension should satisfy $M_{5}^{3}\,V^{\prime\prime}_{\text{gravity}}(d)\propto\delta(d)-(2\pi R)^{-1}$, where $M_{5}^{3}$ is the five-dimensional Planck mass, $d$ is the distance in $y$ between the brane and anti-brane, and the constant term $(2\pi R)^{-1}$ is required for consistency as can be seen by integrating both sides of the equation over the compact $y$-direction. Ultimately, this term should arise from the gravitational back-reaction of the branes. It follows that
@@
where $C_{1}$ an order unity constant. As a cross-check, $V_{\mathrm{gravity}}^{\prime}$ vanishes in the antipodal configuration where $d=\pi R$ as expected due to the symmetry of the system under $d\to-d$. On the other hand, the axion mass gives rise to an effectively repulsive ‘force’ described by the quadratic potential
@@
where $C_{2}$ is another dimensionless constant. Expanding both potentials around the antipodal configuration at $d=\pi R$, we find that the (stable) axion potential dominates over the (tachyonic) gravitational one if $\mathcal{Q}^{2}\gtrsim\mathcal{T}^{\,2}M_{5}^{-3}$, which is independently required by the weak gravity conjecture for 3-brane gauge fields Arkani-Hamed:2006emk ; Ibanez:2015fcv ; Palti:2020mwc ; Kaloper:2023vrs . At the same time, due to Eq. (25), a sub-Planckian axion mass requires $\mathcal{Q}^{2}\ll 2\pi R\,\mathcal{T}\,M_{\mathrm{pl}}^{2}$. Using $M_{\mathrm{pl}}^{2}\sim 2\pi R\,M_{5}^{3}$, we see that both conditions are compatible for $\mathcal{T}\ll M_{\mathrm{pl}}^{4}$. In other words, within our simple EFT it appears possible to stabilize the brane position while keeping the axion light and satisfying the weak gravity conjecture. Whether these competing forces could ultimately lead to a stable configuration within string theory requires a more detailed model building, which we leave for future studies. [^11] In superstring theory, D-branes (of like-charges) are BPS (i.e. the brane charge saturates its mass) objects and exert no force on each other when static, as the NSNS (including gravity) and RR contributions exactly cancel each other Polchinski:1995mt . Applied to our case, the brane and anti-brane would therefore attract each other due to both gravity and gauge contributions to the potential. This might modify the stability condition derive here.
@@ 3.1 Ingredients of M-Theory
where $\mu=0\,,\,\cdots,\,3$ and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the $U(1)$ gauge field strength on the D3-brane. Moreover, $T_{\text{D}3}=T_{\text{M}5}\,\Gamma$ is the D3-brane tension. Here, $A_{\mu}$ originates from the open string modes that we have seen in Eq. (50). These open strings on the D3-brane carry Chan-Paton charges at their ends, and they can move freely within the brane. Note that we have ignored the dependence on the dilaton (which is effectively set to a constant) for simplicity of the discussion. We will stick to this simplification through the rest of the paper.
@@ 3.2 Toroidal Compactification of the M5-Brane
with $\text{M}=0\,,\,\cdots,\,9$ , $f^{9}=y_{0}+R\,\pi(x)$ , $R=\Gamma^{-1}$, and tension $T_{\text{D3}}=T_{\text{M5}}\,\Gamma$ . Note that we have set the fundamental length scale to one here. In this form, the dual action receives an interpretation as a D3-brane in type IIB superstring theory (see discussions around Eq. (53)), with $R$ the radius of the compact direction that is transverse to the D3-brane. In the decompactification limit where $R\rightarrow\infty$ we recover type IIB superstring theory in ten dimensions Schwarz:1995dk ; Aspinwall:1995fw  [^17] A heuristic argument for why compactifying eleven-dimensional M-theory over a two-torus that shrinks to zero leads to ten-dimensional type IIB superstring theory is given below: Take one of the cycles of the two-torus to be the M-theory circle, compactifying on which leads to type IIA superstring theory. Then, we T-dualize the remaining cycle of the two-torus to get type IIB superstring theory. It then follows that shrinking the original cycle in the IIA theory decompactifies the dual circle, which leads to IIB theory in ten dimensions. This D3-brane is localized at $y_{0}$ in the $X^{9}$ direction, which is compactified over a circle of an effective radius $R$ . We observe that the effective radius $R$ of the $X^{9}$ circle is identified with the inverse area of the two-torus, over which the M5-brane is compactified. At the quadratic order in $\pi(x)$ , we recover the free scalar theory (4), where $\pi(x)$ is interpreted as a massless axion field.
@@ 3.3 Three-From Potential from M-Theory
Here, the second term is related to the $\theta$-coupling associated with the QCD instanton,  [^18] In the literature, it is standard to refer to $C^{(0)}$ as an axion Becker:2006dvp . which is distinct from the axion-like particle as a brane-position mode that we are proposing in this paper – the latter is an open string excitation.
@@
Next, we dualize $\mathbb{F}_{\mu\nu\rho}$ in the action. (79), which means forming the ‘parent’ action by adding in the same generating functional (65), such that
@@ Acknowledgements.
We would like to thank Aleksandr Chatrchyan, Georgi Dvali, Nemanja Kaloper, Johannes Lahnsteiner, Niels A. Obers, Antonio Padilla, and Martin S. Sloth for useful discussions. The work of F.N. is supported by VR Starting Grant 2022-03160 of the Swedish Research Council. The work of Z.Y. is supported in part by Olle Engkvists Stiftelse Project Grant 234-0342, VR Project Grant 2021-04013, and the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie Grant Agreement No. 31003710.

### v2

@@ Abstract
Scalar fields with masses protected by global shift symmetries, commonly referred to as axions, are abundantly used in effective field theories in cosmology and particle physics. However, global symmetries cannot be expected to be protected at the fundamental level. Finding consistent ultra-violet completions for axions is therefore a necessity. In this work, we identify the axion with the position mode of a charged 3-brane in (4+1)-dimensions. The shift symmetry of the axion is then a residual diffeomorphism in the fifth dimension orthogonal to the brane. Meanwhile, the brane is coupled to a flux in the fifth dimension. From the (3+1)-dimensional perspective, this construction generates (perturbatively) a mass for the axion and matches previously known proposals in the literature based on the coupling between the axion and a three-form gauge field. In a second step, we uplift this (4+1)-dimensional model to M-theory, where the same three-form is found to couple to the membrane with a (2+1)-dimensional worldvolume. In particular, our proposal also elucidates the duality between the axion and a two-form gauge field in the literature. We show that this dual two-form couples to the boundary of an open membrane in M-theory. Finally, we comment on the relations to and differences from other closed and open string axion monodromy models.
@@ 2.3 Mass Generation from Three-Form Potential
$\int_{\mathcal{M}_{4}}C^{(4)}=\frac{1}{4!}\int_{\mathcal{M}_{4}}\text{d}^{4}x\,\epsilon^{\mu\nu\rho\sigma}\,\partial_{\mu}f^{\text{A}}\,\partial_{\nu}f^{\text{B}}\,\partial_{\rho}f^{\text{C}}\,\partial_{\sigma}f^{\text{D}}\,C_{\mathrm{ABCD}}(f)\,.$ (21)
@@ 2.4 Dimensional Reduction and Anti-Brane
Figure 3: An illustration of the antipodal two-brane configuration. Our universe as a 3-brane sits at $y=0$ and its anti-brane at $y=\pi\,R$ , with $y$ the compactified dimension that is orthogonal to the brane. These two branes carry the same charge $\mathcal{Q}$ but with opposite signs. The background value of the four-form field strength $\bar{H}_{01234}$ is a discontinuous function in $y$, which is equal to $q_{\text{S}}$ in the southern semi-circle and $q_{\text{N}}$ in the northern semi-circle.
Here we provide a more self-contained setting that takes into account the back-reaction of the brane charges on the bulk fluxes. As a result, we will show that the axion potential in the four-dimensional EFT becomes multi-valued, similar to models of axion monodromy. As we will see, the self-consistency of the boundary condition in the compact direction requires that we generalize the action (19) to include an anti 3-brane with opposite charge. For now, we ignore gravitational effects but will get back to the stability of our setting later. We also assume that they have the same tension $\mathcal{T}$ and are placed at $y_{1}=0$ and $y_{2}=R\,\pi$ in the compact dimension. The five-dimensional action (19) is now generalized to be
@@
$\partial_{y}\bar{H}_{01234}=\mathcal{Q}_{1}\,\delta\bigl(y-y_{1}\bigr)+\mathcal{Q}_{2}\,\delta\bigl(y-y_{2}\bigr)\,.$ (34)
Integrating this equation over the circle yields $\mathcal{Q}_{1}+\mathcal{Q}_{2}=0$ , which is what we have imposed. This implies the necessity of introducing the anti-brane. We integrate Eq. (34) to obtain
@@
In $S^{(2)}_{\text{rest}}$ , the first line does not make any contribution to the dynamics of the axions and we thus drop it in the following discussion. In the second line, we encounter a term linear in $a_{1}-a_{2}$ . The coefficient of this latter linear terms is proportional to $\partial_{y}\bar{C}_{0123}=\bar{H}_{01234}$ , which is an undetermined constant on the brane itself. Following the rest of the derivations in Section 2.3 by dualizing $\delta C_{\mu\nu\rho}$ , we find the analog of Eq. (30), with
@@
which further implies that the axion theory (39) admits a discrete gauge symmetry,  [^10] Strictly speaking, this is not a symmetry but instead a transformation between different superselection sectors.
@@
Two massive branes attract each other gravitationally. A fully quantitative analysis would require solving the coupled system involving the five-dimensional gravitational field, the higher-form flux, and a possibly non-trivial dilaton profile, but here we limit ourselves to preliminary remarks. Define $d$ to be the distance in $y$ between the brane and anti-brane. If the extra dimension is stabilized with size $2\pi R$, symmetry dictates that $V_{\mathrm{gravity}}^{\prime}(d)$ vanishes in the antipodal configuration $d=\pi R$, where $d$ is the brane and anti-brane separation along $y$, because neither hemisphere is preferred (see also Fig. 3). Moreover, for $d\ll R$ the compactness of the extra dimension becomes irrelevant and $V_{\mathrm{gravity}}(d)$ should scale linearly with $d$, analogous to the Coulomb potential between two charged plates.  [^11] See related discussions in Burgess:2001fx . A simple gravitational potential that satisfies these requirements alongside the invariance of $V_{\mathrm{gravity}}(d)$ under $d\to-d$ is given by
@@
where $M_{5}^{3}$ is the five-dimensional Planck mass and $C_{1}$ an order unity constant.  [^12] The potential (46) can be thought of as solving the Poisson equation $M_{5}^{3}\,V^{\prime\prime}_{\text{gravity}}(d)\propto\delta(d)-(2\pi R)^{-1}$, where the constant term $(2\pi R)^{-1}$ is required for consistency as can be seen by integrating both sides of the equation over the compact $y$-direction. Mathematically, it is possible to modify the constant to be a periodic function in the Poisson equation, but this would drastically change the background. Another possibility is that one of the branes should have a negative tension.
On the other hand, the axion mass gives rise to an effectively repulsive ‘force’ described by the quadratic potential
@@
where $C_{2}$ is another dimensionless constant. Expanding both potentials around the antipodal configuration at $d=\pi R$, we find that the (stable) axion potential dominates over the (tachyonic) gravitational one if $\mathcal{Q}^{2}\gtrsim\mathcal{T}^{\,2}M_{5}^{-3}$, which is independently required by the weak gravity conjecture for 3-brane gauge fields Arkani-Hamed:2006emk ; Ibanez:2015fcv ; Palti:2020mwc ; Kaloper:2023vrs . At the same time, due to Eq. (25), a sub-Planckian axion mass requires $\mathcal{Q}^{2}\ll 2\pi R\,\mathcal{T}\,M_{\mathrm{pl}}^{2}$. Using $M_{\mathrm{pl}}^{2}\sim 2\pi R\,M_{5}^{3}$, we see that both conditions are compatible for $\mathcal{T}\ll M_{\mathrm{pl}}^{4}$. In other words, within our simple EFT it appears possible to stabilize the brane position while keeping the axion light and satisfying the weak gravity conjecture. Whether these competing forces could ultimately lead to a stable configuration within string theory requires a more detailed model building, which we leave for future studies. [^13] In superstring theory, D-branes (of like-charges) are BPS (i.e. the brane charge saturates its mass) objects and exert no force on each other when static, as the NSNS (including gravity) and RR contributions exactly cancel each other Polchinski:1995mt . Applied to our case, the brane and anti-brane would therefore attract each other due to both gravity and gauge contributions to the potential. This might modify the stability condition derived here in a string theory embedding. In contrast, in the five-dimensional EFT setup, the higher-form gauge field is topological and its effect is captured by the axion due to Eq. (29).
@@ 3.1 Ingredients of M-Theory
where $\mu=0\,,\,\cdots,\,3$ and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the $U(1)$ gauge field strength on the D3-brane. Moreover, $T_{\text{D}3}$ is the D3-brane tension. Here, $A_{\mu}$ originates from the open string modes that we have seen in Eq. (50). These open strings on the D3-brane carry Chan-Paton charges at their ends, and they can move freely within the brane. Note that we have ignored the dependence on the dilaton (which is effectively set to a constant) for simplicity of the discussion. We will stick to this simplification through the rest of the paper.
@@ 3.2 Toroidal Compactification of the M5-Brane
with $\text{M}=0\,,\,\cdots,\,9$ , $f^{9}=y_{0}+R\,\pi(x)$ , $R=\Gamma^{-1}$, and the tension $T_{\text{D3}}=T_{\text{M5}}\,\Gamma$ . Note that we have set the fundamental length scale to one here. In this form, the dual action receives an interpretation as a D3-brane in type IIB superstring theory (see discussions around Eq. (53)), with $R$ the radius of the compact direction that is transverse to the D3-brane. In the decompactification limit where $R\rightarrow\infty$ we recover type IIB superstring theory in ten dimensions Schwarz:1995dk ; Aspinwall:1995fw  [^19] A heuristic argument for why compactifying eleven-dimensional M-theory over a two-torus that shrinks to zero leads to ten-dimensional type IIB superstring theory is given below: Take one of the cycles of the two-torus to be the M-theory circle, compactifying on which leads to type IIA superstring theory. Then, we T-dualize the remaining cycle of the two-torus to get type IIB superstring theory. It then follows that shrinking the original cycle in the IIA theory decompactifies the dual circle, which leads to IIB theory in ten dimensions. This D3-brane is localized at $y_{0}$ in the $X^{9}$ direction, which is compactified over a circle of an effective radius $R$ . We observe that the effective radius $R$ of the $X^{9}$ circle is identified with the inverse area of the two-torus, over which the M5-brane is compactified. At the quadratic order in $\pi(x)$ , we recover the free scalar theory (4), where $\pi(x)$ is interpreted as a massless axion field.
@@ 3.3 Three-From Potential from M-Theory
Note that there is no linear term in $F_{\mu\nu}$ because it is an anti-symmetric tensor. Here, the second term is related to the $\theta$-coupling associated with the QCD instanton,  [^20] In the literature, it is standard to refer to $C^{(0)}$ as an axion Becker:2006dvp . which is distinct from the axion-like particle as a brane-position mode that we are proposing in this paper – the latter is an open string excitation.
@@
Next, we dualize $\mathbb{F}_{\mu\nu\rho}$ in the action. (79), which means constructing the ‘parent’ action by adding in the same generating functional (65), such that
@@ Acknowledgements.
We would like to thank Aleksandr Chatrchyan, Georgi Dvali, Nemanja Kaloper, Johannes Lahnsteiner, Niels A. Obers, Tomás Ortín, Antonio Padilla, and Martin S. Sloth for useful discussions. The work of F.N. is supported by VR Starting Grant 2022-03160 of the Swedish Research Council. The work of Z.Y. is supported in part by Olle Engkvists Stiftelse Project Grant 234-0342, VR Project Grant 2021-04013, the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie Grant Agreement No. 31003710, and the Villum Young Investigator Programme under project No. 71589. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184.